OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 6 x – y = 3 −1 5 −2 2

OpenStudy (anonymous):

The answer should be 2. I used elimination

OpenStudy (anonymous):

2

OpenStudy (anonymous):

I multiplied the second equation by 4 which gave me:4x-4y=12 Then I subtracted to eliminate the x 4x-4x=0x 2y-(-4y)=6y 6-12=-6 6y=-6 y=-1 Now you must plug in the y value into one of the equations and solve for x: x-(-1)=3 x=2